PLL Skip?

tigersfan1

Member
Joined
Sep 27, 2010
Messages
17
I havent had a PLL Skip in a long time like more than 100 solves. I know CFOP and I know the 2look OLL and the one look PLL. Is a reason I havent had a PLL Skip in a long time. Because I use to get it all the time
 
There are 72 ways the last layer can be permuted. There are only 21 algs cause a lot of those cases can appear 4 ways (we just AUF them into a standard position for our algs).
Edit: double ninja'd. Damn Ipod touch pad)
 
I understand now. this is kind off topic but is there any difference between getting a black cube vs. a white cube, does the eye respond differently to different colors?
 
When edges are all oriented. (2Look OLL)

42 Cases for corners to be permuted (inc OLL skip)----> 1/7 of corners being permuted after peforming default alg ------> 1/5 Chance of Edges being skipped because of 5 edge permutation (including solved state)

So if you use 2Look OLL 1/7 x 1/5= 1/35 chance of a PLL skip
Full OLL is a different story though

This is how i think it works with 2Look OLL correct me if im wrong :P
 
When edges are all oriented. (2Look OLL)

42 Cases for corners to be permuted (inc OLL skip)----> 1/7 of corners being permuted after peforming default alg ------> 1/5 Chance of Edges being skipped because of 5 edge permutation (including solved state)

So if you use 2Look OLL 1/7 x 1/5= 1/35 chance of a PLL skip
Full OLL is a different story though

This is how i think it works with 2Look OLL correct me if im wrong :P

Uh, no. PLL isn't affected whatsoever between the use of OCLL or just plain OLL. It all depends on the algs one uses.

PLL breakdown:
1 - solved
1 - H
2 - E
2 - Z
4 - Aa
4 - Ab
4 - F
4 - Ga
4 - Gb
4 - Gc
4 - Gd
4 - J
4 - L
4 - Na
4 - Nb
4 - Ra
4 - Rb
4 - T
4 - Ua
4 - Ub
4 - V
4 - Y
The numbers are the amount of distinctive permutations with their AUFs. Basically, the reason why the solved case has one and a U-perm has four is that if you do a U turn, the solved case looks the same where the U-perm "rotates".
 
When edges are all oriented. (2Look OLL)

42 Cases for corners to be permuted (inc OLL skip)----> 1/7 of corners being permuted after peforming default alg ------> 1/5 Chance of Edges being skipped because of 5 edge permutation (including solved state)

So if you use 2Look OLL 1/7 x 1/5= 1/35 chance of a PLL skip
Full OLL is a different story though

This is how i think it works with 2Look OLL correct me if im wrong :P

All of your numbers are off. The PLL probabilities remain the same whether or not you do Full OLL, as long as you don't do anything that affects the corner or edge permutation (OLLCP, ZBLL etc). The probability of an CPLL skip is 1/6 and the probability of an EPLL skip is 1/12 and 1/6 * 1/12 = 1/72 which gives the probability of a PLL skip, and it'll always be that way if you're not forcing the permutation of the LL during the OLL step(s).


PLL breakdown:
1- solved
1 - H
2 - E
2 - Z
4 - Aa
4 - Ab
4 - F
4 - Ga
4 - Gb
4 - Gc
4 - Gd
4 - J
4 - L
1 - Na
1 - Nb

4 - Ra
4 - Rb
4 - T
4 - Ua
4 - Ub
4 - V
4 - Y
Fixed. 4(1) + 2(2) + 16(4) = 72
 
Last edited:
PLL breakdown:

I'm afraid your numbers dont add up. I think N perms only have 2 each but I cant see what else is wrong at the mo.
Edit: wait no N's only have one each, cause you can do those algs with no AUFs (same as H)
Edit2: f@&£ ninja'd again tonight. I'm going to have to start bumping 3 year old threads so I can be certain I'm saying something new :)
 
Last edited:
Not really, look at feliks he has gotten a couple WR with black, recently he has switched to white for some reason and he got a new WR

He didn't switch. White DaYan products are made with a higher quality plastic. The white has a little less friction. So in this case, there is a minor difference.
 
Back
Top